The standard time for an operation in Work-for-Pay Manufacturing Company is 0.25 hour per unit. There are 8 working hours in a day. The base (day rate) pay for an operator of this operation is $14 per hour. Based on these information and using a 120% incentive pay system, what is the daily pay (in $ per day) for an operator who produced 45 units in 8 working hours?

a. 112.00

b. 157.50

c. 166.60

d. Not enough information given to answer this question
Answer #1

Answer: C (166.60)

Rationale:

Time take to produce 1 unit = 0.25 hours

Number of working hours = 8 hours

Number of units that can be produced in 8 hours = 8/0.25

Number of units that can be produced in 8 hours = 32

But the workers are able to produce 45 units in 8 hours.

Number of extra units produced = 45 - 32 = 13

Time it would have taken to produce extra 13 units = 13 × 0.25 hours

Time it would have taken to produce extra 13 units = 3.25 hours

Therefore, workers will get paid normal wage for 8 hours and will receive incentives for 3.25 hours.

Normal hourly wage = $ 14

Incentives = 120%

Total pay = ($14 × 8) + ($14 × 3.25 × 120%)

Total pay = $ 166.60